A short paper with a few citations

Michael W. Richmond^{1}

School of Physics and Astronomy, Rochester Institute of Technology, Rocheter, NY 14623 Received ; accepted

ABSTRACT

A very short paper with just a few citations, designed to illustrate LaTeX’s handling of a bibliography.

1. Introduction

I wrote a paper a long time ago (Richmond & Knapp 1986). Actually, my advisor, Gillian Knapp, wrote most of it; she was kind enough to permit me to put my name first.

Years later, as part of my graduate work, I helped to automate a telescope. Richmond et al. (1993) describes that project.

REFERENCES

Richmond, M., Treffers, R. R., & Filippenko, A. V. 1993, PASP, 105, 1164 Richmond, M. W., & Knapp, G. R. 1986, AJ, 91, 517

This manuscript was prepared with the AAS L^{A}TEX macros v5.2.